OpenStudy (anonymous):

Use the position function to find the velocity at time t = a (HINT: Velocity is the first derivative of the position function). S(t) = t2 - sin√t, a = 0

OpenStudy (anonymous):

@mathsmind can u help?

OpenStudy (harsimran_hs4):

v = ds/dt can you find v?

OpenStudy (anonymous):

thats the formula?

OpenStudy (harsimran_hs4):

yes you need to differentiate displacement once wrt time to get velocity

OpenStudy (anonymous):

ok

OpenStudy (harsimran_hs4):

find out the derivative and tell

OpenStudy (anonymous):

the derivative of t^2 - sin sqrt(t) right?

OpenStudy (anonymous):

\[V(t)=S'(t)=2t-\cos (\sqrt t)*\frac{ 1 }{ 2 }t^{-\frac{ 1 }{ 2 }} \]\[V(a)=V(0)=2(0)-\cos(\sqrt 0)*\frac{ 1 }{ 2 }(0)^{-\frac{ 1 }{ 2 }}\rightarrow V(0)=0\]

OpenStudy (harsimran_hs4):

@genius12 this is not done you should let @onegirl try out herself

OpenStudy (anonymous):

yea i wanna know how i do it

OpenStudy (anonymous):

Well if you look at the work I did above to get the answer, you should be able to figure it out. If you can't understand any of the steps or how I got to a certain point, ask me.

OpenStudy (anonymous):

okay